OpenStudy (anonymous):

3x-2y=13 x=-3-3y Solve system of Equations using substittion

OpenStudy (anonymous):

3(-3-3y)-2y=13 -9-9y-2y=13 -11y=22 y=-2

OpenStudy (turingtest):

plug that into the second equation to find x